OpenStudy (anonymous):

i need someone to check this for me;; 4|v - 5| = 16 4v - 20 = 16 +20 +20 ------------ 4v = 36 /4 /4 ------------ v = 9 4v - 20 = -16 +20 +20 ------------ 4v = 4 /4 /4 --------- v = 1

OpenStudy (anonymous):

no don't use the distributive propetry

OpenStudy (anonymous):

\[\pm (v-5) = 4\] v - 5 = 4 v = 9 - v + 5 = 4 v = 1

OpenStudy (anonymous):

why is there a + sign in front of (v - 5)

OpenStudy (anonymous):

you have to get the numbers in the absolute value sign alone, you would do that by dividing 4 by 16, so then you'll have |v - 5| = 4 Now what's NOT in the absolute value sign, you will turn into negative and positive to get your two soulitons v -5= -4 add 5 v = 1 v-5=4 add 5 v=9

OpenStudy (anonymous):

that means plus and minus

OpenStudy (anonymous):

i still got the same answer..

OpenStudy (anonymous):

it's plus or minus

OpenStudy (anonymous):

oh cool i didn't know you could do that

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i just did that cause i thought you were suppose to do that.